# spacing=nonfrench

### v2

@@ 3 Design
Figure 2 presents the system overview of LDPKiT, which has four main stages: I. post-processing on noise-injected ${\mathcal{D}_{\rm protected}}$ to form ${\mathcal{D}_{\rm infer}}$ for querying, II. cloud model inference with post-processed privacy-protected data in ${\mathcal{D}_{\rm infer}}$, III. local training with inference results from ${\mathcal{M}_{\rm R}}$ and ${\mathcal{D}_{\rm infer}}$, and IV. execution of users’ private inference queries on the local model with ${\mathcal{D}_{\rm priv}}$. SIDP, on the other hand, does not involve model training. It queries ${\mathcal{M}_{\rm R}}$ with single-layer LDP noised data points and accepts ${\mathcal{M}_{\rm R}}$’s outputs directly. Prior to querying ${\mathcal{M}_{\rm R}}$ for inference, LDPKiT creates ${\mathcal{D}_{\rm protected}}$ based on ${\mathcal{D}_{\rm priv}}$ by adding random noise provable by the Laplacian mechanism (i.e., SIDP). Subsequently, LDPKiT generates ${\mathcal{D}_{\rm cand}}$ from ${\mathcal{D}_{\rm protected}}$ by leveraging the post-processing property of $\epsilon$-LDP [23]. The user has the flexibility to construct ${\mathcal{D}_{\rm infer}}$ for ${\mathcal{M}_{\rm R}}$ inference from ${\mathcal{D}_{\rm cand}}$ using methods such as random selection. For larger datasets, more advanced machine learning techniques, such as active learning [25] and core-set strategies [26], can be employed to optimize the query selection process. Once ${\mathcal{M}_{\rm R}}$ returns inference results, LDPKiT enters the local training stage. We train ${\mathcal{M}_{\rm L}}$ on pairs of the ${\mathcal{D}_{\rm infer}}$ and ${\mathcal{M}_{\rm R}}$’s outputs, aiming to make ${\mathcal{M}_{\rm L}}$ achieve satisfactory accuracy on the original (noise-free) ${\mathcal{D}_{\rm priv}}$. This approach enables LDPKiT to preserve inference utility while safeguarding the privacy of ${\mathcal{D}_{\rm priv}}$ from ${\mathcal{M}_{\rm R}}$. LDPKiT can also be applied to an online learning setting where the user can iterate through the entire process and periodically train ${\mathcal{M}_{\rm L}}$ using ${\mathcal{M}_{\rm R}}$’s predictions on new inference queries.
3.1 Preliminary Experiments
As demonstrated in Section 2.1, model extraction with OOD public data is ineffective. To perform successful model extraction that enables the user to annotate their private data locally on a trustworthy platform, our intuition is to use samples that are in-distribution with ${\mathcal{D}_{\rm priv}}$ to query ${\mathcal{M}_{\rm R}}$, while maintaining privacy through $\epsilon$-LDP. One way is to add $\epsilon$-LDP noise to in-distribution ${\mathcal{D}_{\rm priv}}$ (i.e., ${\mathcal{D}_{\rm protected}}$). However, the limited size of ${\mathcal{D}_{\rm protected}}$ is insufficient for model training. Hence, we need to find a way to generate additional data points to form ${\mathcal{D}_{\rm cand}}$ from ${\mathcal{D}_{\rm priv}}$ that protects privacy in ${\mathcal{D}_{\rm priv}}$ while preserving a distribution as close to ${\mathcal{D}_{\rm priv}}$ as possible.
We test our intuition by systematically introducing random noise to query samples. We employ a two-layer noise injection mechanism. First, we add a base layer of $\epsilon$-LDP noise to each original private data sample, the same way as SIDP. In addition, we apply an additional layer of $\epsilon$-LDP random noise as a post-processing step. Then, we use the post-processed noisy data samples as queries for inference on ${\mathcal{M}_{\rm R}}$. We describe two mechanisms for applying the second layer of $\epsilon$-LDP noise in the next section. Further, we train the ${\mathcal{M}_{\rm L}}$ model on the post-processed noisy samples labelled by the ${\mathcal{M}_{\rm R}}$. This approach enhances the ${\mathcal{M}_{\rm L}}$’s prediction accuracy beyond that of the ${\mathcal{M}_{\rm R}}$ over repeated queries while ensuring consistent privacy guarantees. The base noise remains intact throughout the querying process, and the post-processing of the noised private samples complies with the post-processing property of LDP. A detailed proof of the privacy guarantees is provided in Appendix A.
To realize our idea, we conduct a proof-of-concept experiment on Fashion-MNIST. Specifically, we construct a relatively small ${\mathcal{D}_{\rm priv}}$ with 500 samples randomly selected from Fashion-MNIST. We set ${\mathcal{M}_{\rm R}}$ as ResNet-152 and ${\mathcal{M}_{\rm L}}$ as ResNet-18. To ensure robust and reliable results, we average the accuracies over three runs and across three different dataset splits, mitigating the influence of randomness. We calibrate the $\epsilon$ value to 2.0 so that ${\mathcal{M}_{\rm R}}$’s inference accuracy under SIDP (with a single layer of noise applied to each data point) is only a few multiples of the probability of random guessing (e.g., 15% to 30%). This configuration ensures that ${\mathcal{M}_{\rm R}}$ provides minimal yet sufficient information while preserving privacy. The recorded SIDP accuracy is 29.86%. We then implement and evaluate our above-mentioned two-layer noise injection mechanism. We post-process 500 SIDP queries with an extra layer of random noise to form an inference dataset, ${\mathcal{D}_{\rm infer}}$ with 249,500 data points (i.e., its maximum available dataset size with 499 randomly noised variations of the SIDP noisy queries). We then train ${\mathcal{M}_{\rm L}}$ with ${\mathcal{D}_{\rm infer}}$ and ${\mathcal{M}_{\rm R}}$’s noisy predictions on those noisy data. The trained ${\mathcal{M}_{\rm L}}$’s inference accuracy on the 500 original (noise-free) ${\mathcal{D}_{\rm priv}}$ samples is improved to 43.56% (i.e., 13% accuracy increase compared to SIDP) without compromising the privacy level. We collect the $p$ value results from the dependent two-sample t-test for statistical significance and the value is negligible; hence, the improvement is solid. However, the utility recovery achieved through this mechanism does not yield an optimal trade-off between utility and privacy. One reason is that ${\mathcal{D}_{\rm priv}}$ is too small (i.e., 500 data samples). More importantly, the idea of post-processing the noised private data with an extra layer of random noise is intuitive yet undirected. Thus, we further investigate whether more advanced strategies can be devised to design the post-processing LDP noise applied to ${\mathcal{D}_{\rm priv}}$ in forming ${\mathcal{D}_{\rm infer}}$, enabling the ${\mathcal{M}_{\rm L}}$ to learn more meaningful information about the decision boundaries and achieve improved accuracy. We elaborate on our design of LDP noise in Section 3.2 and discuss the results in Section 4.
3.2 LDPKiT’s Noise Injection Mechanism
@@
LDPKiT applies Laplacian noise to the privacy-sensitive queries in ${\mathcal{D}_{\rm priv}}$ to generate ${\mathcal{D}_{\rm infer}}$ for cloud model inference, ensuring the $\epsilon$-LDP guarantee [22]. We define $\epsilon$ in the context of LDP; therefore, the same amount of noise is added to each data sample in ${\mathcal{D}_{\rm priv}}$ based on the $\epsilon$ value. We assume that queries in ${\mathcal{D}_{\rm priv}}$ are independent and identically distributed (i.i.d.), and the formulation of $\epsilon$-LDP is valid for each individual data point in ${\mathcal{D}_{\rm priv}}$. Consequently, the privacy leakage is non-cumulative and bounded by $\epsilon$ per query. Given a private dataset ${\mathcal{D}_{\rm priv}}$ containing $|{\mathcal{D}_{\rm priv}}|$ i.i.d. data points, our goal is to create a ${\mathcal{D}_{\rm infer}}$ dataset for querying ${\mathcal{M}_{\rm R}}$ and training ${\mathcal{M}_{\rm L}}$ with privacy safeguarded by $\epsilon$-LDP. To achieve this, we employ a two-layer noise injection framework. We first add Laplacian noise with a scale inversely proportional to $\epsilon$ to each data point $d_{\text{priv}_{i}}\in{\mathcal{D}_{\rm priv}}$ to form ${\mathcal{D}_{\rm protected}}$ with the same size as ${\mathcal{D}_{\rm priv}}$, where $i=1,2,\dots,|{\mathcal{D}_{\rm priv}}|$. This first layer of base Laplacian noise remains intact throughout the entire post-processing and inference process for privacy protection. Then, unlike the conventional use of differentially private Laplacian noise (i.e., SIDP), LDPKiT applies the second layer of noise to ${\mathcal{D}_{\rm protected}}$ adhering to the post-processing property in two alternative ways. This layer of noise aims to expand the ${\mathcal{D}_{\rm infer}}$ for knowledge transfer and recovering the labels of ${\mathcal{D}_{\rm priv}}$. Details are described as follows.
Addition of an extra layer of random noise that satisfies ${\epsilon}$-LDP (LDPKiT-Rand). As a baseline approach, we apply post-processing to each data point in ${\mathcal{D}_{\rm protected}}$ by adding an additional layer of Laplacian noise ($L$) with the same scale in the form of $d_{\text{cand}_{i}}\in{\mathcal{D}_{\rm cand}}=d_{\text{priv}_{i}}\in{\mathcal{D}_{\rm priv}}+L_{1}+L_{2}$ for $i=1,2,\dots,|{\mathcal{D}_{\rm priv}}|$. Although theoretically, we can generate an infinite number of random versions of each ${\mathcal{D}_{\rm protected}}$ data point to construct ${\mathcal{D}_{\rm cand}}$, we cap the final dataset ${\mathcal{D}_{\rm cand}}$ to have at most $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$ data points, aligning with the second mechanism in the next section for evaluation purposes.
@@
Superimposition of other noised data points in ${\mathcal{D}_{\rm infer}}$ that also satisfies ${\epsilon}$-LDP (LDPKiT-Sup). LDPKiT-Rand introduces unrelated noise that may cause the resulting data to deviate further from the distribution of ${\mathcal{D}_{\rm priv}}$. To address this and generate more in-distribution samples, we adopt a different approach designed to better preserve the characteristics of ${\mathcal{D}_{\rm priv}}$. After applying the first layer of $\epsilon$-LDP random noise to each data point in ${\mathcal{D}_{\rm priv}}$, we exhaustively combine these modified data points in all possible pairs with averaged pixel-wise addition. The candidate data point $d_{\text{cand}_{i}}\in{\mathcal{D}_{\rm cand}}$ is computed as $d_{\text{cand}_{i}}=(d_{\text{protected}_{i}}+d_{\text{protected}_{j}})/2,i,j=1,2,\dots,|{\mathcal{D}_{\rm priv}}|$. This process results in at most $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$ distinct permutations in the protected candidate dataset ${\mathcal{D}_{\rm cand}}$.
@@
When querying ${\mathcal{M}_{\rm R}}$ with ${\mathcal{D}_{\rm infer}}$, the user has the option to utilize either the entire available dataset (${\mathcal{D}_{\rm cand}}$) or a randomly selected subset of it. We analyze the effect of the size of ${\mathcal{D}_{\rm infer}}$ ($|{\mathcal{D}_{\rm infer}}|$) in Section 4.4. In Appendix A, we prove that the post-processed noisy data points in the final inference dataset ${\mathcal{D}_{\rm infer}}$, generated using both LDPKiT-Rand and LDPKiT-Sup, adhere to $\epsilon$-LDP and satisfy the post-processing property of LDP.
@@
Figure 3: Sample of a noised Fashion-MNIST data point with label 5 (sandal) and $\epsilon$ set to 1.5.
4 Evaluation
@@
In this section, we discuss the evaluation results and answer our research questions in Section 1 with empirical analysis.
4.1 Experimental setup
We evaluate LDPKiT on three diverse datasets, namely SVHN [27], Fashion-MNIST [28] and PathMNIST from MedMNIST2D for medical imaging in pathology [29]. For the ML models, we use ResNet-152 [30] as ${{\mathcal{M}_{\rm R}}}$, and ResNet-18 [30] and MobileNetV2 [31] as ${{\mathcal{M}_{\rm L}}}$. We assume that ${\mathcal{M}_{\rm R}}$ only returns the hard labels. The ${\mathcal{M}_{\rm L}}$ models are initialized with random weights. As a reference, when no privacy protection exists and thus no noise is added, ${\mathcal{M}_{\rm R}}$’s average accuracies on ${\mathcal{D}_{\rm priv}}$ of Fashion-MNIST, SVHN and PathMNIST are 93.33%, 94.73% and 86.2%, respectively. The accuracies on an unseen validation dataset, ${\mathcal{D}_{\rm val}}$, are 93.22%, 96.30% and 81.24%, respectively.
We construct different ${\mathcal{M}_{\rm R}}$ for different tasks and datasets. To allocate data points for ${\mathcal{D}_{\rm priv}}$, instead of using the default training split, we train ResNet-152 (${\mathcal{M}_{\rm R}}$) on 35k, 48,257 and 89,996 data points for Fashion-MNIST, SVHN and PathMNIST, respectively. For Fashion-MNIST, the candidate pool of ${\mathcal{D}_{\rm priv}}$ has 25k data points with ${\mathcal{D}_{\rm val}}$ comprising 10k. For SVHN, the candidate pool for ${\mathcal{D}_{\rm priv}}$ is 25k and ${\mathcal{D}_{\rm val}}$ has 26,032 data points. For PathMNIST, the candidate pool for ${\mathcal{D}_{\rm priv}}$ has 10,004 data points and ${\mathcal{D}_{\rm val}}$ has 7,180. To simulate a scenario aligned with our setting,, where users have limited private data requiring label information, we set $|{\mathcal{D}_{\rm priv}}|=1,500$ for experiments in Sections 4.2 and 4.5. Specifically, we construct ${\mathcal{D}_{\rm priv}}$ by randomly selecting 1,500 balanced data points from the candidate pool to mitigate potential uncertainties associated with class imbalances. The effect of varying sizes of ${\mathcal{D}_{\rm infer}}$ and ${\mathcal{D}_{\rm priv}}$ is studied in Section 4.4. All experiments are repeated over three random seeds on three random subset splits to determine the statistical significance of our findings. The results of the dependent two-sample t-test confirm that all improvements are statistically significant (i.e., $p<0.05$).
We record SIDP accuracy based on the labels returned by ${{\mathcal{M}_{\rm R}}}$ for ${\mathcal{D}_{\rm protected}}$, whereas LDPKiT accuracy is determined based on labels from ${\mathcal{M}_{\rm L}}$ for ${\mathcal{D}_{\rm priv}}$ where ${\mathcal{M}_{\rm L}}$ is trained on ${\mathcal{M}_{\rm R}}$’s labels on ${\mathcal{D}_{\rm infer}}$. To calibrate $\epsilon$ values per dataset, we ensure that the SIDP accuracy of ${\mathcal{M}_{\rm R}}$ corresponds to $1.5-3X$ random guessing (e.g., 15% to 30% accuracy for a 10-class dataset), allowing ${\mathcal{M}_{\rm R}}$ to provide minimal but meaningful information for training ${\mathcal{M}_{\rm L}}$. Specifically, we use $\epsilon$ values of 2.0, 1.5, and 1.25 for SVHN, 2.0 and 1.5 for Fashion-MNIST, and 10.0 and 7.0 for PathMNIST. Figure 3 shows a privacy-protected data point in Fashion-MNIST’s ${\mathcal{D}_{\rm cand}}$ generated with LDPKiT-Rand and LDPKiT-Sup. Additional noisy samples are included in Appendix C. Details of our hyperparameter choices are discussed and reported in Appendix B.
@@ 4.1 Experimental setup
Figure 4: Accuracy comparisons on ${\mathcal{D}_{\rm priv}}$ with calibrated $\epsilon$ values: SIDP versus ResNet-18(${\mathcal{M}_{\rm L}}$) trained using LDPKiT-Rand and LDPKiT-Sup.
(a) SVHN
(b) Fashion-MNIST
(c) PathMNIST

### v3

@@ III Design
Figure 2 describes both SIDP and LDPKiT. LDPKiT has four main stages: I. post-processing on noise-injected $\mathcal{D}_{\rm protected}$} to form a larger inference dataset, $\mathcal{D}_{\rm infer}$}, II. remote model inference with post-processed privacy-protected data from $\mathcal{D}_{\rm infer}$}, III. local training with inference results from $\mathcal{M}_{\rm R}$} and $\mathcal{D}_{\rm infer}$}, and IV. execution of users’ private inference queries on $\mathcal{M}_{\rm L}$} with $\mathcal{D}_{\rm priv}$}. $\mathcal{D}_{\rm infer}$} is generated from $\mathcal{D}_{\rm protected}$} using LDPKiT, with its size controlled by a user-defined query budget, as each query to $\mathcal{M}_{\rm R}$} incurs a financial cost. To optimize query efficiency, methods such as active learning [34] and core-set selection [35] could be used to choose informative subsets from $\mathcal{D}_{\rm infer}$}. We leave the investigation of such optimization techniques to future work and, for now, we assume that the user randomly samples a subset from the maximum available $\mathcal{D}_{\rm infer}$} for querying. LDPKiT may also be applied to an online learning setting where the user can iterate through the entire process and periodically train $\mathcal{M}_{\rm L}$} using $\mathcal{M}_{\rm R}$}’s predictions on new inference queries. SIDP, in contrast, does not involve $\mathcal{M}_{\rm L}$} training. It queries $\mathcal{M}_{\rm R}$} directly with data points from $\mathcal{D}_{\rm protected}$} and accepts $\mathcal{M}_{\rm R}$}’s outputs directly. Since $|{\mathcal{D}_{\rm protected}}|<|{\mathcal{D}_{\rm infer}}|$, SIDP has a lower querying cost than LDPKiT, but at the cost of utility.
III-A Preliminary Experiments
[…]
To validate our idea, we conduct a preliminary experiment on SVHN. Specifically, we construct a relatively small $\mathcal{D}_{\rm priv}$} with 500 samples randomly selected from SVHN. We set ResNet-152 as $\mathcal{M}_{\rm R}$} and ResNet-18 as $\mathcal{M}_{\rm L}$}. To ensure reliable results, we average the accuracies over three runs and across three different dataset splits. We set $\epsilon=1.5$ so that $\mathcal{M}_{\rm R}$}’s inference accuracy under SIDP (with a single layer of noise applied to each data point) is only 14.16%. This configuration ensures that $\mathcal{M}_{\rm R}$} provides minimal yet sufficient information while preserving privacy. To evaluate the effectiveness of our two-layer data augmentation mechanism, we then apply an additional layer of random noise 499 times to each of the 500 samples in $\mathcal{D}_{\rm protected}$}, creating an augmented inference dataset, $\mathcal{D}_{\rm infer}$}, with 249,500 data points (i.e., $n=499$). Performing model extraction with $\mathcal{D}_{\rm infer}$} results in an $\mathcal{M}_{\rm L}$} with 34.07% accuracy, while performing extraction only with the 500 original samples from $\mathcal{D}_{\rm protected}$} results in an $\mathcal{M}_{\rm L}$} that essentially guesses randomly (i.e., 10% accuracy). Compared to querying directly using $\mathcal{D}_{\rm protected}$} (i.e., SIDP), which gets 14.16% accuracy, this provides a 20% increase with no loss of privacy. A dependent two-sample t-test result shows that the improvement is statistically significant ($p<0.05$).
However, this naïve random-noise-based utility recovery mechanism does not offer an optimal utility-privacy trade-off, not only due to the limited size of $\mathcal{D}_{\rm priv}$} (500 samples) but also the undirected nature of the random noise augmentation. We hypothesize that data distributions form structured latent spaces, where queries constructed closer to decision boundaries could provide more informative training signals. Motivated by this insight, we explore a superimposition-based augmentation approach that refines the post-processed LDP noise. This approach aims to help $\mathcal{M}_{\rm L}$} better capture the decision boundaries, ultimately achieving higher accuracy on $\mathcal{D}_{\rm priv}$}. Under the same experimental setup, ResNet-18 achieves 45.69% accuracy—a 10% improvement over the naïve random noise augmentation.
III-B LDPKiT’s Data Augmentation Mechanism
@@
LDPKiT applies Laplacian noise to the privacy-sensitive data in $\mathcal{D}_{\rm priv}$} to generate $\mathcal{D}_{\rm infer}$} for remote model inference, ensuring the $\epsilon$-LDP guarantee [28]. Since $\epsilon$ is defined in the context of LDP, an equal amount of noise is added to each data sample in $\mathcal{D}_{\rm priv}$} based on the $\epsilon$ value. We assume that queries in $\mathcal{D}_{\rm priv}$} are independent and identically distributed (i.i.d.), and the formulation of $\epsilon$-LDP is valid for each individual data point in $\mathcal{D}_{\rm priv}$}. Consequently, the privacy leakage is non-cumulative and bounded by $\epsilon$ per query. Given a private dataset $\mathcal{D}_{\rm priv}$} containing $|{\mathcal{D}_{\rm priv}}|$ i.i.d. data points, our goal is to create an augmented privacy-preserving $\mathcal{D}_{\rm infer}$} dataset safeguarded by $\epsilon$-LDP for querying $\mathcal{M}_{\rm R}$} and training $\mathcal{M}_{\rm L}$}. We propose two methods for data augmentation: the first involves adding different random noise samples to each data sample in $\mathcal{D}_{\rm protected}$} as described in Section III-A. The second involves superimposing pairs of samples from $\mathcal{D}_{\rm protected}$}. We emphasize in both cases that since the data augmentation is performed on $\mathcal{D}_{\rm protected}$}, the privacy level of both augmentation techniques remains the same as the original $\mathcal{D}_{\rm protected}$} under the post-processing property of LDP.
LDPKiT-Rand. We formally describe the augmentation technique mentioned previously in Section III-A here: for each $d_{\text{priv}_{i}}\in{\mathcal{D}_{\rm priv}}$, we construct an augmented dataset of the form: $d_{\text{infer}_{i}}\in{\mathcal{D}_{\rm infer}}=d_{\text{priv}_{i}}+\mathcal{L}_{i}+\mathcal{L}_{j}$, where $\mathcal{L}_{i}$ and $\mathcal{L}_{j}$ are Laplacian noise samples drawn independently with the same scale. The first noise layer, $\mathcal{L}_{i}$, is used to generate the initial privacy-protected dataset: $d_{\text{protected}_{i}}=d_{\text{priv}_{i}}+\mathcal{L}_{i}$, where $i=1,2,\dots,|{\mathcal{D}_{\rm protected}}|$. We then apply the second noise layer, $\mathcal{L}_{j}$, where $j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|-1$, as post-processing to create an expanded augmented dataset for querying. While it is theoretically possible to generate an infinite number of such noisy variants per data point, we cap the size of the maximum possible augmented dataset $\mathcal{D}_{\rm infer}$} to $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ to make it comparable to LDPKiT-Sup below.
@@
LDPKiT-Sup. LDPKiT-Rand introduces unrelated noise that may cause the resulting data to deviate further from the distribution of $\mathcal{D}_{\rm priv}$}. LDPKiT-Sup is thus proposed as an alternative to better preserve the characteristics of $\mathcal{D}_{\rm priv}$}. After applying the first layer of $\epsilon$-LDP random noise to each data point in $\mathcal{D}_{\rm priv}$} to produce $\mathcal{D}_{\rm protected}$}, we combine pairwise samples of $\mathcal{D}_{\rm protected}$} for all pairs by averaging corresponding pixel values. Each augmented inference data point $d_{\text{infer}_{i}}\in{\mathcal{D}_{\rm infer}}$ is computed as $d_{\text{infer}_{i}}=avg(d_{\text{protected}_{i}},d_{\text{protected}_{j}})\text{ where }i,j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|,i\neq j$. This process results in at most $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ permutations in the protected inference dataset $\mathcal{D}_{\rm infer}$}.
@@
Post-processed data samples in the augmented inference dataset $\mathcal{D}_{\rm infer}$}, generated with LDPKiT-Rand and LDPKiT-Sup, adhere to $\epsilon$-LDP and satisfy the post-processing property of LDP in Definition II.3. When querying $\mathcal{M}_{\rm R}$} with $\mathcal{D}_{\rm infer}$}, the user has the option to utilize either the entire available augmented inference dataset or a subset of it. We analyze the effect of $|{\mathcal{D}_{\rm infer}}|$ in Section IV-D. In Appendix A-A, we prove that the post-processed noisy data points in the augmented inference dataset $\mathcal{D}_{\rm infer}$}, generated using both LDPKiT-Rand and LDPKiT-Sup, adhere to $\epsilon$-LDP and satisfy the post-processing property of LDP.
@@
Figure 3: Example of a noised Fashion-MNIST data point with label 5 (sandal) and $\epsilon$ set to 1.5.
IV Evaluation
@@
In this section, we discuss the evaluation results and answer our research questions in Section I with empirical analysis.
@@
Figure 4: Inference accuracy comparisons on $\mathcal{D}_{\rm priv}$} with various $\epsilon$: ResNet-152 ($\mathcal{M}_{\rm R}$})’s SIDP versus $\mathcal{M}_{\rm L}$} trained using LDPKiT-Rand and LDPKiT-Sup. The results show the rank of performance as LDPKiT-Sup > LDPKiT-Rand > SIDP
IV-A Experimental setup
We evaluate LDPKiT on three diverse datasets: SVHN [36], Fashion-MNIST [37] and PathMNIST from MedMNIST2D for medical imaging in pathology [38]. For ML models, we use ResNet-152 [39] as $\mathcal{M}_{\rm R}$}}, and ResNet-18 [39] and MobileNetV2 [40] as $\mathcal{M}_{\rm L}$}}. The $\mathcal{M}_{\rm L}$} models are initialized with random weights.
To allocate data points for $\mathcal{D}_{\rm priv}$}, instead of using the default training split, we train ResNet-152 ($\mathcal{M}_{\rm R}$}) on 35k, 48,257 and 89,996 data points for Fashion-MNIST, SVHN and PathMNIST, respectively. A holdout dataset, $\mathcal{D}_{\rm val}$}, is used to evaluate $\mathcal{M}_{\rm L}$}’s generalizability. For Fashion-MNIST, the candidate pool of $\mathcal{D}_{\rm priv}$} has 25k data points and $\mathcal{D}_{\rm val}$} has 10k. For SVHN, the candidate pool for $\mathcal{D}_{\rm priv}$} has 25k data points and $\mathcal{D}_{\rm val}$} has 26,032. For PathMNIST, the candidate pool for $\mathcal{D}_{\rm priv}$} has 10,004 data points and $\mathcal{D}_{\rm val}$} has 7,180. To simulate a scenario that aligns with our setting, where users have limited private data, we set $|{\mathcal{D}_{\rm priv}}|=1,500$ for experiments in Sections IV-B. Specifically, we construct $\mathcal{D}_{\rm priv}$} by randomly selecting 1,500 data points in a balanced manner from the candidate pool so that each class is equally represented (balanced). This balanced sampling helps approximate identical distribution conditions across classes, reducing potential biases introduced by class imbalance. The effect of varying sizes of $\mathcal{D}_{\rm infer}$} and $\mathcal{D}_{\rm priv}$} is studied in Section IV-D. All experiments are repeated over three random seeds on three random subset splits to determine the statistical significance of our findings. The results of the dependent two-sample t-test confirm that all improvements are statistically significant (i.e., $p<0.05$).
We report the baseline SIDP accuracy using the labels returned by $\mathcal{M}_{\rm R}$}} on $\mathcal{D}_{\rm protected}$}, while the LDPKiT accuracy reflects the accuracy of $\mathcal{M}_{\rm L}$} on $\mathcal{D}_{\rm priv}$}, where $\mathcal{M}_{\rm L}$} is trained using $\mathcal{M}_{\rm R}$}}’s labels on $\mathcal{D}_{\rm infer}$}. We follow the industry standards [29, 30] to set $\epsilon$ for each dataset and ensure that the SIDP accuracy of $\mathcal{M}_{\rm R}$} corresponds to $1.5-3\times$ random guessing (e.g., 15% to 30% accuracy for a 10-class dataset), allowing $\mathcal{M}_{\rm R}$} to provide minimal but meaningful information for training $\mathcal{M}_{\rm L}$}. Moreover, we make sure that the noise added is high enough so that LDPKiT-protected data is resilient to data reconstruction attacks (See Section IV-E). Specifically, we choose $\epsilon$ values of 2.0, 1.5, and 1.25 for SVHN, 2.0 and 1.5 for Fashion-MNIST, and 10.0 and 7.0 for PathMNIST. The $\epsilon$ choices are consistent among experiments of SIDP, LDPKiT-Rand and LDPKiT-Sup. Figure 3 shows a privacy-protected data sample in Fashion-MNIST’s $\mathcal{D}_{\rm infer}$} generated with LDPKiT-Sup and LDPKiT-Rand. Additional noisy samples are shown in Appendix A-B. Hyperparameter choices are discussed and reported in Appendix A-C.